ln(x-3)+ln(x)=ln(4)
Answer by venugopalramana(3286)

ln(x-3)+ln(x)=ln(4)
USE LN X+LN Y=LN XY
LN X(X-3)=LN 4
TAKING ANTILOGS
X(X-3)=4
X^2-3X-4=0
X^2-4X+X-4=0
X(X-4)+1(X-4)=0
(X-4)(X+1)=0
X=4 OR -1...SINCE X=-1 GIVES L0G OF NEGATIVE NUMBERS WE DISCARD THAT (LOG OF A -VE NUMBER DOES NOT EXIST)SO X=4